Q? using only five odd numbers, add them up so that the result is 32, the five odd numbers have to be whole numbers only.
Answer by venugopalramana(3286)

Now coming to this specific problem
5 odd numbers can never add up to an even
number .it is impossible.that is why ,i thought there
might be some error in your question and gave the
answer assuming 6 numbers.you can understand this
taking 1,2,3,4,5,6..etc odd numbers add them and
check.mathematically..an odd number is given by
2n+1..so 5 times that is 5(2n+10)=10n+5..10n being
always even ,when 5 is added to it it will become odd
and never even .
it is only by guess you can do taking first 5 consecutive positive odd numbers to get a sum of above 32 we get 3+5+7+9+11=35..so we take -1+1+5+7+9+11=32..as the answer